Consider the spot curve for hypothetical annual-coupon U.S. Treasury securities given in the table below (only the first 5 years of the curve are provided). Calculate the missing one-year forward rates f(T*,1) indicated by “????” and add them to the table. Show your work for all calculations in the space below the table.

Year 1.0 2.0 Spot Rate (%) 5.50 6.02 6.55 6.87 7.20 Forward Rate (%) 5.50 ???? ???? ???? ???? 3.0 4.0 5.0
Answer #1

1-year Forward rate would be:

F[1,1] = (1 + 6.02% 2 -= 6.54% (1 + 5.5%)

(1 + 6.55%) F[2, 1) = (1 +6.02%)2 = 7.62%

(1 + 6.87% F[3, 1) = (1 + 6.559 13 = 7.84%

(1 +7.20%) F[4, 1] = (1 + 6.87%) 4 = 8.53%

Year Spot rate Forward rate
1 5.50% 5.50%
2 6.02% 6.54%
3 6.55% 7.62%
4 6.87% 7.84%
5 7.20% 8.53%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.